### 09 Apr 2026 v1

# Thermal Time and Irreversibility from Non-Commuting Observables in Accelerated Quantum Systems

Abstract
We investigate the operational meaning of temporal ordering in relativistic quantum field theory using localized detector models. The presence of a time parameter in the dynamical description does not by itself guarantee that different sequences of operations correspond to physically distinguishable processes. We show that such distinguishability arises when two conditions are simultaneously satisfied: the underlying quantum state exhibits Kubo–Martin–Schwinger (KMS) structure, and the detector couples through non-commuting observables.
We analyze uniformly accelerated two-level detectors interacting with a quantum field in the Minkowski vacuum. The restriction of the vacuum to the detector trajectory induces a thermal response characterized by the Unruh temperature and, equivalently, by a Tolman redshift profile along the trajectory. For detectors coupled sequentially through distinct internal observables, the reduced detector state depends on the ordering of the interactions. This dependence appears already at second order in perturbation theory and is controlled by the same parameter that governs the KMS condition.
The ordering asymmetry is first detected at the level of the reduced density matrices, which differ for distinct interaction protocols. Its informational content is then quantified using quantum relative entropy. In a minimal two-level model, the relevant states form a family of non-commuting Gibbs states with identical spectra and different generators. The relative entropy between these states admits a closed-form expression and depends only on the dimensionless parameter set by the local temperature and the detector energy scale.
We further analyze the corresponding information geometry by comparing the Bogoliubov–Kubo–Mori and Bures metrics. Their difference in non-commuting directions provides a quantitative distinction between irreversible entropic cost and operational distinguishability. Both quantities are governed by the same thermal scale fixed by acceleration.
These results provide an operational probe of thermal time as the modular flow generated by $K=-\log\rho$, with normalization set by the local temperature. Irreversibility arises from the mismatch between inequivalent modular structures associated with non-commuting observables, and is quantified by relative entropy. In this framework, temporal ordering becomes physically meaningful when it can be encoded in the state of a system and extracted through non-commuting measurements.
Independent researcher
marcello@gravity.phys.nagoya-u.ac.jp
Keywords thermal time $\cdot$ Unruh effect $\cdot$ KMS states $\cdot$ quantum information geometry $\cdot$ irreversibility
1 Introduction
The emergence of irreversible behavior from time-reversal invariant dynamics remains a central issue in the foundations of statistical mechanics and quantum theory. Traditional explanations rely on coarse-graining, special initial conditions, or typicality arguments, but these approaches do not directly address when a temporal ordering parameter acquires observable significance at the level of microscopic processes.
In relativistic quantum field theory, the question becomes sharper. The notion of time depends on the choice of observer, and even the classification of states as thermal or non-thermal may vary between observers. A key example is the Unruh effect, according to which a uniformly accelerated observer perceives the Minkowski vacuum as a thermal state [1, 2]. This phenomenon is naturally formulated using the Kubo–Martin–Schwinger (KMS) condition, which characterizes equilibrium states in quantum statistical mechanics [4, 5]. A detailed review is given in [3].
From an algebraic perspective, the Unruh effect is closely related to the Bisognano–Wichmann theorem, which identifies the modular flow of wedge algebras with Lorentz boosts [6, 7]. This result provides a concrete realization of modular structure in quantum field theory and underlies the thermal time hypothesis proposed by Connes and Rovelli [8]. In that framework, a state determines a canonical flow, suggesting that temporal structure is state-dependent rather than fundamental.
Detector models provide an operational framework for probing these ideas. The Unruh–DeWitt detector [9, 10] describes a localized quantum system interacting with a field along a worldline. Its response is determined by the field’s correlation functions evaluated along that trajectory. The theory of such detectors has been developed in detail, including careful treatments of switching functions and response rates [11, 12, 13].
In standard analyses, detectors couple to the field through a single observable, and the focus is on excitation probabilities and thermal spectra. This approach does not address how the internal algebra of detector observables affects the extraction of information from the field. In particular, it does not consider whether different sequences of interactions can be distinguished at the level of the detector state.
The present work focuses on this question. We consider detectors that couple sequentially through distinct internal observables that do not commute. This introduces a potential dependence on the ordering of interactions. The central result is that such ordering dependence becomes physically meaningful when the field state satisfies a KMS condition. In this case, the correlation functions distinguish forward and reversed sequences through a thermally weighted structure.
This effect admits a natural interpretation in terms of quantum information geometry. When restricted to commuting deformations of a Gibbs state, monotone metrics such as the Bures and Bogoliubov–Kubo–Mori (BKM) metrics coincide. For non-commuting deformations, they differ. The Bures metric captures operational distinguishability, while the BKM metric arises from the second-order expansion of relative entropy [14]. This distinction allows one to quantify the difference between distinguishability and irreversible entropy in a precise manner.
In the accelerated setting, the relevant dimensionless parameter is fixed by the Unruh temperature and depends on the detector’s position in Rindler space. For a minimal two-level system, the relative entropy between detector states associated with different internal observables can be computed exactly. This yields an explicit relation between acceleration, distinguishability, and irreversible entropy.
The analysis establishes a connection between three structures: KMS thermality, non-commutativity of observables, and relative entropy. Together, these determine when temporal ordering becomes observable in detector-based measurements.
The paper is organized as follows. Section 2 introduces the algebraic and detector-theoretic framework. Section 3 defines the detector model with non-commuting couplings. Section 4 analyzes the ordering dependence of the reduced detector state in accelerated motion. Section 5 provides the information-geometric interpretation in terms of relative entropy and modular structure. Section 6 summarizes the results and discusses possible extensions.
2 Algebraic and physical framework
We collect the elements required for the analysis, focusing on the structures that directly enter the detector dynamics.
Let $\mathcal{A}$ be the algebra of observables of a quantum field and let $\omega$ be a state on $\mathcal{A}$. For a scalar field $\phi$, the two-point correlation function is given by
$W(x,x^{\prime})=\omega(\phi(x)\phi(x^{\prime})).$
When evaluated along a worldline $x(\tau)$, this defines a function $W(\tau,\tau^{\prime})$ that governs the response of localized detectors. In perturbation theory, detector transition probabilities and reduced states depend only on this two-point function [10, 11, 13].
A state is said to satisfy the KMS condition at inverse temperature $\beta$ with respect to a flow parameter $\tau$ if correlation functions admit an analytic continuation and satisfy
$\omega(A(\tau)B(\tau^{\prime}))=\omega(B(\tau^{\prime})A(\tau+i\beta))$
for suitable observables $A$ and $B$ [4, 5]. This condition provides a representation-independent characterization of thermal equilibrium.
A physically relevant realization arises for uniformly accelerated observers. For a trajectory with constant proper acceleration $a$, the restriction of the Minkowski vacuum to the corresponding worldline satisfies the KMS condition with inverse temperature
$\beta=\frac{2\pi}{a}.$
This is the Unruh effect [1, 2], which can be understood algebraically through the Bisognano–Wichmann theorem [6, 7].
We model the detector as a two-level system interacting with the field through an Unruh–DeWitt coupling. The interaction Hamiltonian in the interaction picture is
$H_{I}(\tau)=\lambda\,\chi(\tau)\,\mu(\tau)\,\phi(x(\tau)),$
where $\lambda$ is a coupling constant, $\chi(\tau)$ is a switching function, and $\mu(\tau)$ acts on the detector Hilbert space. The corresponding evolution operator is
$U=\mathcal{T}\exp\left(-i\int d\tau\,H_{I}(\tau)\right),$
with $\mathcal{T}$ denoting time ordering. For an initial product state $\rho_{D}\otimes\rho_{\phi}$, the reduced detector state after the interaction is
$\rho_{D}^{\prime}=\mathrm{Tr}_{\phi}\left(U(\rho_{D}\otimes\rho_{\phi})U^{\dagger}\right).$
Expanding perturbatively, the leading contributions depend on integrals of the Wightman function along the trajectory [10, 11, 13].
The framework described above provides the connection between field correlations and detector states. In the following sections, this structure is extended by allowing the detector to couple through multiple internal observables, thereby introducing non-commutativity at the level of the detector algebra.
3 Detector model with non-commuting couplings
We now introduce the detector setup used to probe the interplay between non-commutativity and the correlation structure of the quantum field.
We consider a two-level detector with Hilbert space $\mathcal{H}_{D}\simeq\mathbb{C}^{2}$, moving along a prescribed worldline $x(\tau)$. The detector interacts with a scalar field through couplings that act on its internal degrees of freedom. In contrast with the standard formulation, we allow the detector to couple through distinct internal observables.
Specifically, we introduce two Hermitian operators
$\mu_{x}=\sigma_{x},\qquad\mu_{y}=\sigma_{y},$
where $\sigma_{x}$ and $\sigma_{y}$ are Pauli matrices satisfying $\sigma_{x},\sigma_{y}]=2i\sigma_{z}$. These operators define two independent interaction channels.
We consider interaction Hamiltonians of the form
$H_{x}(\tau)=\lambda\,\chi_{x}(\tau)\,\sigma_{x}\,\phi(x(\tau)),\qquad H_{y}(\tau)=\lambda\,\chi_{y}(\tau)\,\sigma_{y}\,\phi(x(\tau)),$
where $\chi_{x}(\tau)$ and $\chi_{y}(\tau)$ are smooth switching functions with compact, non-overlapping support. This ensures that the detector interacts with the field through $\sigma_{x}$ and $\sigma_{y}$ in a well-defined sequence. Similar perturbative constructions with localized switching are standard in detector theory [11, 12, 13].
Because the supports do not overlap, the time-ordered evolution operator factorizes as
$U_{x\to y}=U_{y}U_{x},\qquad U_{y\to x}=U_{x}U_{y},$
where
$U_{x}=\mathcal{T}\exp\left(-i\int d\tau\,H_{x}(\tau)\right),\qquad U_{y}=\mathcal{T}\exp\left(-i\int d\tau\,H_{y}(\tau)\right).$
We compare the reduced detector states obtained from the two protocols. For an initial state $\rho_{D}\otimes\rho_{\phi}$, we define
$\rho_{x\to y}=\mathrm{Tr}_{\phi}\left(U_{y}U_{x}(\rho_{D}\otimes\rho_{\phi})U_{x}^{\dagger}U_{y}^{\dagger}\right),$
$\rho_{y\to x}=\mathrm{Tr}_{\phi}\left(U_{x}U_{y}(\rho_{D}\otimes\rho_{\phi})U_{y}^{\dagger}U_{x}^{\dagger}\right).$
The difference between these two states defines the ordering-dependent contribution to the detector dynamics,
$\Delta\rho_{D}=\rho_{x\to y}-\rho_{y\to x}.$
A nonvanishing $\Delta\rho_{D}$ indicates that the two interaction protocols lead to distinct physical outcomes. The ordering of the interactions is thus encoded in the reduced detector state.
To evaluate this quantity, we expand the evolution operators perturbatively in the coupling $\lambda$ and isolate the cross terms associated with the two interaction channels. The ordering dependence arises at second order and can be expressed in terms of the commutator of the integrated interaction operators.
A detailed derivation is provided in Appendix B. The resulting ordering-dependent contribution takes the form
$\Delta\rho_{D}=i\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,G^{(1)}(\tau,\tau^{\prime})\,[\sigma_{z},\rho_{D}]+O(\lambda^{3}),$
where
$G^{(1)}(\tau,\tau^{\prime})=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau)$
is the Hadamard function along the detector trajectory.
This expression shows that the ordering asymmetry is governed by the interplay between the non-commutativity of the detector couplings and the symmetrized correlation structure of the field.
This result makes the structure of the asymmetry explicit. The dependence on the ordering of the interactions is controlled by two ingredients: the non-commutativity of the detector couplings, encoded in $\sigma_{x},\sigma_{y}]$, and the state-dependent symmetrized field correlation function $G^{(1)}$. If either ingredient is absent, the ordering-dependent contribution vanishes.
The quantity $\Delta\rho_{D}$ establishes that the ordering of non-commuting interactions is encoded in the reduced detector state. At this stage, the result is purely dynamical: it shows that different protocols lead to different states, but does not yet quantify their distinguishability or thermodynamic significance. This will be addressed in terms of relative entropy and information geometry in the following sections.
4 Accelerated motion and KMS-induced asymmetry
We now evaluate the ordering-dependent contribution obtained in Section 3 for uniformly accelerated motion. The relevant quantity is the Hadamard function
$G^{(1)}(\tau,\tau^{\prime})=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau),$
whose pullback to the accelerated trajectory inherits a thermal structure from the KMS property of the vacuum state restricted to the Rindler wedge.
Let the detector follow a trajectory with constant proper acceleration $a$. The restriction of the Minkowski vacuum to this trajectory defines a KMS state at inverse temperature
$\beta=\frac{2\pi}{a}.$
Equivalently, for the family of coaccelerated Rindler observers one may write the local Tolman profile as
$T(\xi)\sqrt{-g_{00}(\xi)}=\text{const},$
and, for the standard Rindler metric, this yields the local inverse temperature
$\beta(\xi)=2\pi\xi$
in natural units. The KMS condition implies
$W(\tau,\tau^{\prime})=W(\tau^{\prime},\tau+i\beta),$
together with the usual analyticity properties [4, 5, 3].
For a stationary trajectory, the correlation functions depend only on the time difference, so it is convenient to introduce the Fourier transforms
$W(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{W}(\omega)e^{-i\omega(\tau-\tau^{\prime})},$
$G^{(1)}(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)e^{-i\omega(\tau-\tau^{\prime})}.$
The KMS condition implies the detailed-balance relation
$\widetilde{W}(-\omega)=e^{-\beta\omega}\widetilde{W}(\omega),$
and therefore
$\widetilde{G}^{(1)}(\omega)=\widetilde{W}(\omega)+\widetilde{W}(-\omega)=\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega).$
If one introduces the spectral function
$\widetilde{\Delta}(\omega)=\widetilde{W}(\omega)-\widetilde{W}(-\omega),$
this may also be written as
$\widetilde{G}^{(1)}(\omega)=\coth\!\left(\frac{\beta\omega}{2}\right)\widetilde{\Delta}(\omega).$
This is the thermal weighting relevant to the ordering asymmetry derived in Section 3.
Substituting the Fourier representation of $G^{(1)}$ into the exact second-order expression for $\Delta\rho_{D}$, one obtains
$\Delta\rho_{D}=i\lambda^{2}[\sigma_{z},\rho_{D}]\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)\,\widetilde{\chi}_{x}(-\omega)\widetilde{\chi}_{y}(\omega)+O(\lambda^{3}),$
where
$\widetilde{\chi}_{j}(\omega)=\int_{-\infty}^{\infty}d\tau\,\chi_{j}(\tau)e^{i\omega\tau}$
denotes the Fourier transform of the switching function.
Using the KMS relation, this becomes
$\Delta\rho_{D}=i\lambda^{2}[\sigma_{z},\rho_{D}]\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega)\,\widetilde{\chi}_{x}(-\omega)\widetilde{\chi}_{y}(\omega)+O(\lambda^{3}).$
The ordering asymmetry is therefore controlled by the same local inverse temperature that determines the Unruh response and, equivalently, by the Tolman redshift profile along the accelerated trajectory.
This formula shows that the role of the KMS condition is not merely to make the detector thermal in the usual excitation-rate sense. It fixes a universal relation between the positive- and negative-frequency components of the field correlations and hence a universal thermal weighting of the symmetrized correlator that governs the ordering dependence of the reduced state.
The result also clarifies the inertial case. Along an inertial trajectory in the Minkowski vacuum, the Wightman function does not satisfy a KMS condition with respect to the detector proper time. The Hadamard function is still nonvanishing, so ordering-dependent contributions may arise from the combination of non-commutativity and field correlations. However, there is no universal detailed-balance relation and hence no thermal factor that organizes the asymmetry. Uniform acceleration singles out a regime in which the ordering dependence is governed by a well-defined local temperature.
The quantity $\Delta\rho_{D}$ therefore detects a genuine dynamical asymmetry between the two protocols, while the KMS condition endows this asymmetry with a universal thermal structure. In the next section, we turn to the question of how this state difference should be quantified informationally and how it is related to modular flow and thermal time.
5 Relative entropy, information geometry, and thermal time
The ordering asymmetry derived in the previous section admits a quantitative interpretation in terms of relative entropy and information geometry. This section has two purposes. The first is to identify the appropriate state-space quantities that measure the asymmetry detected by $\Delta\rho_{D}$. The second is to make explicit the meaning of thermal time in the present setting.
We consider the effective equilibrium states associated with detectors in a uniformly accelerated setting. The relevant temperature is not a fixed external parameter but the local temperature associated with the family of coaccelerated observers. In Rindler space this local temperature satisfies the Tolman law
$T(\xi)\sqrt{-g_{00}(\xi)}=\text{const},$
and for the uniformly accelerated trajectory considered here this reduces to the Unruh profile
$\beta(\xi)=2\pi\xi.$
The thermal state seen by the detector is therefore position-dependent even though the ambient state is the Minkowski vacuum.
We focus on the minimal case in which the detector couples through two distinct internal observables. Let
$H_{x}=\Delta\sigma_{x},\qquad H_{y}=\Delta\sigma_{y},$
where $\Delta$ sets the detector energy scale. The corresponding local Gibbs states are
$\rho_{x}=\frac{e^{-\beta H_{x}}}{Z},\qquad\rho_{y}=\frac{e^{-\beta H_{y}}}{Z},$
with
$Z=\mathrm{Tr}(e^{-\beta H_{x}})=\mathrm{Tr}(e^{-\beta H_{y}}).$
Introducing the dimensionless parameter
$s=\beta\Delta,$
these states take the explicit form
$\rho_{x}=\frac{e^{-s\sigma_{x}}}{2\cosh s},\qquad\rho_{y}=\frac{e^{-s\sigma_{y}}}{2\cosh s}.$
The two states have identical spectra but different generators. This is the simplest non-commuting Gibbs family relevant to the detector problem. It is also the point at which thermal time becomes nontrivial. For each Gibbs state, the modular generator is
$K_{x}=-\log\rho_{x}=\beta H_{x}+\log Z,\qquad K_{y}=-\log\rho_{y}=\beta H_{y}+\log Z.$
Up to the additive scalar term $\log Z$, which does not affect the modular automorphism group, the two generators are
$K_{x}\sim\beta H_{x},\qquad K_{y}\sim\beta H_{y}.$
Thermal time is therefore the flow generated by the local modular Hamiltonian, and in the present setting it depends explicitly on the local temperature through $\beta(\xi)$. The Tolman–Unruh profile determines how that modular scale varies along the accelerated trajectory.
The operator difference between the two detector states is not, by itself, an informational quantity. It detects that the two states are distinct, but it is not a canonical informational measure. The natural entropic measure is the quantum relative entropy
$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}(\rho_{y}\log\rho_{y}-\rho_{y}\log\rho_{x}).$
This quantity measures the irreversible informational cost of identifying $\rho_{y}$ with $\rho_{x}$ and is therefore the appropriate object for discussing irreversibility.
Using the identities
$\log\rho_{x}=-\log(2\cosh s)\,I-s\sigma_{x},\qquad\log\rho_{y}=-\log(2\cosh s)\,I-s\sigma_{y},$
one obtains
$D(\rho_{y}\|\rho_{x})=s\,\mathrm{Tr}\big(\rho_{y}(\sigma_{x}-\sigma_{y})\big).$
The expectation value of Pauli operators in a Gibbs state of this form satisfies
$\mathrm{Tr}(\rho_{n}\sigma)=-\tanh s\,n,$
where $n$ is the unit vector specifying the direction of the generator. It follows that
$D(\rho_{y}\|\rho_{x})=s\tanh s.$
This expression provides a closed-form measure of the mismatch between the two Gibbs structures associated with the non-commuting detector couplings. It depends only on the dimensionless parameter $s=\beta\Delta$, which combines the detector energy scale with the local inverse temperature determined by the Unruh effect or, equivalently, by the Tolman profile along the accelerated trajectory.
It is important to distinguish this quantity from the perturbative state difference $\Delta\rho_{D}$ derived in Section 4. The latter establishes that different interaction orderings lead to different reduced detector states and therefore detects the presence of a dynamical asymmetry. However, $\Delta\rho_{D}$ does not by itself provide a measure of distinguishability or irreversibility.
The relative entropy $D(\rho_{y}\|\rho_{x})$ instead quantifies the distinguishability between the effective Gibbs states associated with the two non-commuting couplings. In this sense, it measures the informational content of the ordering asymmetry once expressed in terms of thermal structures.
We now relate this result to information geometry. Let $\rho_{\theta}$ denote the Gibbs state generated by
$H_{\theta}=\Delta(\cos\theta\,\sigma_{x}+\sin\theta\,\sigma_{y}),$
so that $\theta$ parametrizes rotations between the two non-commuting generators. For nearby states, the relative entropy admits a quadratic expansion
$D(\rho_{\theta}\|\rho_{0})\simeq\frac{1}{2}g^{\mathrm{BKM}}_{\theta\theta}\theta^{2},$
where $g^{\mathrm{BKM}}$ is the Bogoliubov–Kubo–Mori metric. For the present family of states, one finds
$g^{\mathrm{BKM}}_{\theta\theta}=s\tanh s.$
By contrast, the Bures metric, which quantifies distinguishability through fidelity, yields
$g^{\mathrm{Bures}}_{\theta\theta}=\tanh^{2}s.$
The ratio between the two metrics is therefore
$\frac{g^{\mathrm{BKM}}_{\theta\theta}}{g^{\mathrm{Bures}}_{\theta\theta}}=\frac{s}{\tanh s}.$
This ratio measures the discrepancy between entropic cost and operational distinguishability for non-commuting deformations of the Gibbs state.
The dependence on acceleration can be made explicit by expressing $s$ in terms of the Rindler coordinate $\xi$. For a uniformly accelerated observer, the local inverse temperature is $\beta=2\pi\xi$, so that
$s=2\pi\xi\Delta.$
In the limit $\xi\to 0$, corresponding to large acceleration,
$D(\rho_{y}\|\rho_{x})\sim s^{2},\qquad\frac{g^{\mathrm{BKM}}}{g^{\mathrm{Bures}}}\to 1.$
In the opposite limit $\xi\to\infty$, corresponding to small acceleration,
$D(\rho_{y}\|\rho_{x})\sim s,\qquad\frac{g^{\mathrm{BKM}}}{g^{\mathrm{Bures}}}\sim s.$
Finally, we connect these results to modular structure. Each Gibbs state defines a modular generator
$K_{x}=-\log\rho_{x}=\beta H_{x}+\log Z,\qquad K_{y}=-\log\rho_{y}=\beta H_{y}+\log Z.$
Up to the additive scalar term $\log Z$, the corresponding modular flows are generated by $\beta H_{x}$ and $\beta H_{y}$. The normalization of these generators is fixed by the local inverse temperature $\beta(\xi)$, reflecting the Tolman relation and the Unruh effect.
The relative entropy $D(\rho_{y}\|\rho_{x})$ therefore measures the mismatch between the modular structures associated with the two non-commuting couplings. In this sense, thermal time is realized as the modular flow selected by the state, while irreversibility is quantified by the relative entropy between inequivalent Gibbs structures.
The detector dynamics derived in Section 4 provides the operational input: the KMS condition fixes the analytic structure of correlation functions, and non-commuting observables probe this structure through sequential interactions. The resulting ordering asymmetry is encoded in the detector state, while its informational content is captured by the relative entropy between the corresponding Gibbs states.
6 Discussion
We have identified a minimal mechanism through which temporal ordering becomes operationally meaningful in relativistic quantum systems. The key ingredients are the KMS structure of the underlying state and the non-commutativity of the observables through which the detector couples to the field.
At the dynamical level, the ordering of interactions is encoded in the reduced detector state. Distinct interaction protocols lead to different final states whenever non-commutativity produces a nonvanishing contribution $\Delta\rho_{D}$. In the presence of a KMS state, this asymmetry acquires a universal thermal structure controlled by the local temperature. This establishes that temporal ordering can be reflected directly in the state of a localized quantum system.
The quantitative content of this asymmetry is captured by relative entropy. In the minimal two-level model, the relevant states form a family of Gibbs states with identical spectra and different generators. The relative entropy between them depends only on the dimensionless parameter $s=\beta\Delta$, linking the strength of the asymmetry to the local temperature scale.
The comparison between the Bogoliubov–Kubo–Mori and Bures metrics clarifies the geometric structure underlying this result. In commuting directions the two metrics coincide, while in non-commuting directions they differ. This difference quantifies the separation between irreversible entropic cost and operational distinguishability, and is governed by the same thermal parameter.
Thermal time is realized in this framework through the modular generator $K=-\log\rho$ associated with a Gibbs state. In the accelerated setting, this generator is normalized by the local inverse temperature $\beta(\xi)$, determined by the Tolman relation and the Unruh effect. The dependence of the relative entropy and of the information-geometric quantities on $s=\beta\Delta$ makes explicit how irreversibility is controlled by this local thermal scale.
The detector model provides a concrete probe of these structures. The KMS condition fixes the analytic properties of the correlation functions, while non-commuting observables probe this structure through sequential interactions. The resulting effective Gibbs states encode the ordering dependence, and their relative entropy quantifies its irreversible content.
The framework suggests several directions for further investigation. One extension is to consider detector systems with higher-dimensional internal Hilbert spaces, where the structure of non-commuting generators is richer. Another is to analyze other spacetime settings in which KMS states arise, such as stationary spacetimes with horizons, where nontrivial temperature profiles may lead to additional effects.
The results show that temporal ordering acquires physical meaning when it is encoded in the state of a system through the interplay of thermality and non-commutativity. In this setting, irreversibility is quantified by relative entropy, and its magnitude is set by the same thermal scale that governs the underlying KMS structure.
Appendix A Relative entropy for non-commuting Gibbs states
We derive the closed-form expression for the relative entropy between two Gibbs states generated by non-commuting operators. This computation provides the entropic measure used in the main text to quantify the mismatch between the corresponding non-commuting Gibbs structures.
We consider a two-level system with density matrices
$\rho_{x}=\frac{e^{-s\sigma_{x}}}{2\cosh s},\qquad\rho_{y}=\frac{e^{-s\sigma_{y}}}{2\cosh s},$
where $s=\beta\Delta$.
Using the identity
$e^{-s\sigma_{n}}=\cosh s\,I-\sinh s\,\sigma_{n},$
one obtains the Bloch representation
$\rho_{n}=\frac{1}{2}\left(I-\tanh s\,\sigma_{n}\right).$
The logarithm of the density matrix follows directly from its Gibbs form,
$\log\rho_{n}=-\log(2\cosh s)\,I-s\sigma_{n}.$
The quantum relative entropy is defined as
$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}(\rho_{y}\log\rho_{y})-\mathrm{Tr}(\rho_{y}\log\rho_{x}).$
Substituting the expressions above, one finds
$D(\rho_{y}\|\rho_{x})=\mathrm{Tr}\left[\rho_{y}\left(-s\sigma_{y}+s\sigma_{x}\right)\right].$
The expectation values of Pauli operators in the state $\rho_{y}$ are
$\mathrm{Tr}(\rho_{y}\sigma_{y})=-\tanh s,\qquad\mathrm{Tr}(\rho_{y}\sigma_{x})=0.$
It follows that
$D(\rho_{y}\|\rho_{x})=s\tanh s.$
More generally, if one interpolates between the two generators by
$H_{\theta}=\Delta(\cos\theta\,\sigma_{x}+\sin\theta\,\sigma_{y}),$
the second-order expansion of relative entropy with respect to $\theta$ yields the Bogoliubov–Kubo–Mori metric used in Section 5.
The expression for $D(\rho_{y}\|\rho_{x})$ depends only on the dimensionless parameter $s=\beta\Delta$ and therefore encodes the dependence of the irreversible informational cost on the local temperature scale. In the context of the main text, this quantity provides the entropic measure of the mismatch between Gibbs states generated by non-commuting observables.
Appendix B Perturbative derivation of the ordering asymmetry
We derive the second-order ordering-dependent contribution to the reduced detector state for the two sequential interaction protocols introduced in Section 3.
Let
$A_{x}=\int d\tau\,H_{x}(\tau),\qquad A_{y}=\int d\tau\,H_{y}(\tau),$
and
$B_{x}=\int d\tau d\tau^{\prime}\,\mathcal{T}\big(H_{x}(\tau)H_{x}(\tau^{\prime})\big),\qquad B_{y}=\int d\tau d\tau^{\prime}\,\mathcal{T}\big(H_{y}(\tau)H_{y}(\tau^{\prime})\big).$
To second order in the coupling, the individual evolution operators are
$U_{x}\approx 1-iA_{x}-\frac{1}{2}B_{x},\qquad U_{y}\approx 1-iA_{y}-\frac{1}{2}B_{y}.$
Since the supports of the switching functions do not overlap, the two ordered protocols are
$U_{x\to y}\approx 1-i(A_{x}+A_{y})-\frac{1}{2}(B_{x}+B_{y})-A_{y}A_{x},$
$U_{y\to x}\approx 1-i(A_{x}+A_{y})-\frac{1}{2}(B_{x}+B_{y})-A_{x}A_{y}.$
For an initial product state $\rho_{D}\otimes\rho_{\phi}$, the reduced detector states are
$\rho_{x\to y}=\mathrm{Tr}_{\phi}\!\left(U_{x\to y}(\rho_{D}\otimes\rho_{\phi})U_{x\to y}^{\dagger}\right),$
$\rho_{y\to x}=\mathrm{Tr}_{\phi}\!\left(U_{y\to x}(\rho_{D}\otimes\rho_{\phi})U_{y\to x}^{\dagger}\right).$
At order $\lambda^{2}$, the first-order contributions and the quadratic sandwich term are identical for both protocols. The difference therefore arises entirely from the ordered cross terms and can be written as
$\Delta\rho_{D}=\rho_{x\to y}-\rho_{y\to x}=\mathrm{Tr}_{\phi}\!\left(\bigl[[A_{x},A_{y}],\,\rho_{D}\otimes\rho_{\phi}\bigr]\right)+O(\lambda^{3}).$
Using
$H_{x}(\tau)=\lambda\,\chi_{x}(\tau)\,\sigma_{x}\,\phi(x(\tau)),\qquad H_{y}(\tau^{\prime})=\lambda\,\chi_{y}(\tau^{\prime})\,\sigma_{y}\,\phi(x(\tau^{\prime})),$
we compute the commutator
$A_{x},A_{y}]=\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,[\sigma_{x}\phi(x(\tau)),\,\sigma_{y}\phi(x(\tau^{\prime}))].$
Using
$A\otimes B,\,C\otimes D]=[A,C]\otimes BD+CA\otimes[B,D],$
together with
$\sigma_{x},\sigma_{y}]=2i\sigma_{z},\qquad\sigma_{y}\sigma_{x}=-i\sigma_{z},$
one finds
$\sigma_{x}\phi(x(\tau)),\,\sigma_{y}\phi(x(\tau^{\prime}))]=i\sigma_{z}\,\{\phi(x(\tau)),\phi(x(\tau^{\prime}))\}.$
Therefore,
$A_{x},A_{y}]=i\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,\sigma_{z}\,\{\phi(x(\tau)),\phi(x(\tau^{\prime}))\}.$
Tracing over the field degrees of freedom gives
$\Delta\rho_{D}=i\lambda^{2}\int d\tau d\tau^{\prime}\,\chi_{x}(\tau)\chi_{y}(\tau^{\prime})\,G^{(1)}(\tau,\tau^{\prime})\,[\sigma_{z},\rho_{D}]+O(\lambda^{3}),$
where
$G^{(1)}(\tau,\tau^{\prime})=\omega\!\left(\{\phi(x(\tau)),\phi(x(\tau^{\prime}))\}\right)=W(\tau,\tau^{\prime})+W(\tau^{\prime},\tau)$
is the Hadamard function.
For stationary trajectories, one may introduce the Fourier transform
$G^{(1)}(\tau,\tau^{\prime})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\widetilde{G}^{(1)}(\omega)e^{-i\omega(\tau-\tau^{\prime})}.$
If the state satisfies the KMS condition,
$\widetilde{W}(-\omega)=e^{-\beta\omega}\widetilde{W}(\omega),$
then
$\widetilde{G}^{(1)}(\omega)=\widetilde{W}(\omega)+\widetilde{W}(-\omega)=\bigl(1+e^{-\beta\omega}\bigr)\widetilde{W}(\omega)=\coth\!\left(\frac{\beta\omega}{2}\right)\widetilde{\Delta}(\omega),$
where $\widetilde{\Delta}(\omega)=\widetilde{W}(\omega)-\widetilde{W}(-\omega)$.
This shows explicitly how the KMS condition induces a thermal weighting of the symmetrized correlator that governs the ordering asymmetry.